Shift Summary 8/29/99 EVENING shift

Hogan Nguyen (hogann@fnal.gov)
Mon, 30 Aug 1999 00:08:57 -0500

On Shift: Ricardo Barbosa, Hogan

Hours of beam available at all intensities: 8.3 h
Hours of beam available at nominal intensity: 8.1 h
Hours of E832 data (special runs not included): 8.0
Hours of special runs (laser scans, muon runs...): 0h

Accounting of lost time: Occasion stream 5 problems from
the PAC10 crate (crate 53)
and the laser crate. Occasional DAQDEAD errors.

========================================================================
Runs ended this shift:

14373 - 35K vac 2pi0 tapes full

Tapes brought to FCC this shift: QK(S,C,K)351-353, QKB122
Accesses during shift:

========================================================================
Overview of shift: fairly smooth shift

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14373 15:30 We inherit smooth run from the day shift

18:10 Dave Northacker installs 3 new ArEth bottles on the west side

20:15 Process CheckSDAQ_from_wisc dies. We restart it with Theo's
help

20:30 Mounted split tape QKN031 onto ktev1-tps3d2

303 20:45 NM1U and NM1H tripped. The beam profile doesn't even show up
on SEED2. We notify MCR who said that it was due to water problem.
We put in the beam stop.

313 21:00 NM1U and NM1H back up. Beam stops out. Everything back to normal
in spill 313.

21:02 Mounted split tape QKNL39 onto ktev2-tps3d2

22:09 Tapes full. Stopped run 14373. Loaded new XT map generated
from run 14372. This should bring sod means back to nominal. The
sods have been slowly drifting up, and just touched the warning
limits.

14374 1 22:26 Run starts. New XT file puts smiles back onto dc and T3 kumacs.

The pmt gain drift plot for the previous run was rather wide,
where as it is narrow for this run. However, I did not update
the CsI constants at the start of this run.

22:40 Several problems with crate 53 and the laser crate. Ultimately
fixed by recycling the laser crate repeater.

23:15 Disk usage >90% warning on ktev1g,2g,4g,3c. These are due to
onsplit files. Sydney cleans up the areas. Back to normal.

========================================================================
Follow up Notes on Previous Shift:

o kept eyes on channel 450. It was very stable during the
time we were.

========================================================================
Items to Follow up on Next Shift:

o if we lose beam for >= 40 minutes, and the DA works, do a laser scan.

o if no beam for > 2 hours, page Julie for divider swap.

o keep your eyes on channel 450 (see log book p. 103). If it remains stable,
and we do a laser calibration, maybe we don't need to swap the divider
(ask Ed).

========================================================================